# Nonlinear Fore(Back)casting and Innovation Filtering for Causal-Noncausal VAR Models

### v3

@@ 5.2 Uncertainty of the Estimated Prediction Set
where $P_{0}$ is the true distribution function of process ($Y_{t}$) and $Q(y,\alpha,P_{0})$ denotes the $\alpha$-quantile of $Y_{1,T+1}$ conditional on $Y_{T}=y$ and derived from the joint multivariate predictive density $l(y|\underline{Y}_{T})$ (see Proposition 1 for the closed-form expression of the predictive density). Let $Q_{l}$ and $Q_{u}$ denote the true $\alpha_{1}/2$ and $21-\alpha_{1}/$ 2 conditional quantiles, respectively. Recall that under the semi-parametric approach, the true distribution $P_{0}$ is characterized by $\Phi_{0}$ and $g_{0}$. By using the expression of the prediction interval in a Gaussian framework, the asymptotically valid prediction interval for $Y_{1,T+1}$ can be equivalently written as:
@@
with  [^10] Note that $\Phi^{-1}(\alpha_{1}/2)$ is negative.
$m(y,\alpha_{1};P_{0})=0.5[Q_{l}(y,\alpha_{1}P_{0})+Q_{u}(y,\alpha_{1};P_{0})],$ (5.3)
$\mbox{and}\;\;\sigma(y,\alpha_{1};P_{0})=-[1/(2\Phi^{-1}(\alpha_{1}/2)][Q_{u}(y,\alpha_{1};P_{0})-Q_{l}(y,\alpha_{1};P_{0})].$ (5.4)
@@
This normalized representation of the prediction interval resembling the traditional Gaussian approach can be used even if the conditional density function of $(Y_{t})$ is not Gaussian. In particular, the functions $y\rightarrow m(y,\alpha_{1};P_{0}),y\rightarrow\sigma(y,\alpha_{1};P_{0})$ are nonlinear, in general.
ii) Estimated Prediction Interval
@@
In our framework, the unknown marginal predictive density function $P_{0}$ can be consistently estimated from its expression (3.12) as $\hat{P}$, given the estimated matrix of autoregressive parameters $\hat{\Phi}$ and the residuals $\hat{\varepsilon}_{t}$ obtained from the semi-parametric GCov estimator that provides the nonparametric estimator $\hat{g}$. Then, we can compute an estimated prediction interval of $Y_{1,T+1}$ :
$\widehat{PI}(y,\alpha_{1})=[m(y,\alpha_{1};\hat{P})\pm\Phi^{-1}(\alpha_{1}/2)\,\sigma(y,\alpha_{1};\hat{P})]=[Q_{l}(y,\alpha_{1};\hat{P}),Q_{u}(y,\alpha_{1};\hat{P})].$ (5.5)
where $Q_{l}(y,\alpha_{1};\hat{P})$ and $Q_{u}(y,\alpha_{1};\hat{P})$ are the $\alpha_{1}/2$ and $21-\alpha_{1}/$ 2 conditional quantiles of the estimated predictive density.
This estimated prediction interval (5.5) is consistent of the true prediction interval (5.2) when the number of observations tends to infinity. We need to distinguish:
a) the true prediction interval $PI(y,\alpha_{1})$ of $Y_{1,T+1}$ satisfying $P_{0}[Y_{1,T+1}\in PI(y,\alpha_{1})|Y_{T}=y]=1-\alpha_{1},\;\;\forall y.$
By construction, the true prediction interval has the correct conditional coverage probability of $1-\alpha_{1}$. It depends on the unknown true distribution, or equivalently on $\Phi_{0},g_{0}$.
b) The estimated prediction interval $\widehat{PI}(y,\alpha_{1})$ that does not satisfy the conditional coverage condition in finite sample: $P_{0}[Y_{1,T+1}\in\widehat{PI}(y,\alpha_{1})|Y_{T}=y]\neq 1-\alpha_{1},\;\;\forall y.$
due to the estimation errors.
iii) Bootstrap Adjusted Prediction Interval
@@
Since this estimated prediction interval is random, its asymptotic distribution can be approximated by bootstrap, which is applied by replicating the trajectory of the process by backcasting. More precisely, given $\hat{\Phi}$, $\hat{g}$ considered fixed and the residuals, we can generate by backcasting the artificial paths $Y_{t}^{s},\;t=1,...,T$ with the same terminal condition $Y_{T}^{s}=Y_{T}=y$ for all the bootstrapped samples (see, Corollary 2 for the closed-form expression of the backward predictive distribution). The backcasting can be performed as the following sequence of one-step backcasts:
@@
Step 2: By replicating the backcasted path $S$ times, we end up generating $S$ bootstrapped series $Y_{t}^{s},t=1,...,T$ of length $T$ equal to the length of the initial series and with the same terminal value $Y_{T}$.
step 3: From each replicated path $(Y_{t}^{s},t=1,...,T)$, we estimate the model parameters $\Phi^{s}$ and $\hat{g}^{s}$, $s=1,...,S$. That allows us for computing at $Y_{T}$ the predictive density estimator $\hat{P}^{s},s=1,...,S$ of $Y_{T+1}$ given $Y_{T}$ and $S$ new prediction intervals of $Y_{1,T+1}$ from each of the replicated paths.
@@
where $\hat{P}^{s}$ is the semi-parametric estimate of $P_{0}$ from the artificial path $(Y_{t}^{s},t=1,...,T)$. Given $\hat{P}^{s},s=1,...,S$ of $Y_{T+1}$ this bootstrap PI of $Y_{1,T+1}$ given $Y_{T}$ can be replicated independently $s=1,...,S$ times.
The components of the prediction interval (5.6) are denoted by:
@@
step 5: For large $S$, the joint sample distribution of $[\hat{m}^{s}(y,\alpha_{1}),\hat{\sigma}^{s}(y,\alpha_{1})]$ provides an approximation of the distribution of $[m(y,\hat{P}),\sigma(y,\hat{P})]$, when $T$ is finite and sufficiently large.
5.3 Confidence Interval of the Prediction Interval
The estimated prediction interval $\widehat{PI}(y,\alpha_{1})$ is a pointwise estimator of interval $PI(y,\alpha_{1})$ and as any estimator, is random itself. This randomness is difficult to assess since $\widehat{PI}(y,\alpha_{1})$ is a random set (interval). Let us now extend the method of pointwise estimation to build confidence sets for $PI(y,\alpha_{1})$. Since there does not exist a total ordering on intervals, we first constrain the confidence set to be of the form:
$\widehat{PI}(y,\alpha_{1},q)=[m(y,\alpha_{1};\hat{P})\pm q\,\sigma(y,\alpha_{1};\hat{P})].$ (5.8)
and search for an estimator of $q$ providing the correct asymptotic coverage. This confidence set has the following coverage probability of the true prediction interval:
$\Pi_{0}(y,\alpha_{1};q)=P_{0}[\widehat{PI}(y,\alpha_{1},q)\supset PI(y,\alpha_{1})|Y_{T}=y]$ (5.9)
$=$ $P_{0}[m(y,\alpha_{1};\hat{P})-q\,\sigma(y,\alpha_{1};\hat{P})<m(y,\alpha_{1};P_{0})+\Phi^{-1}(\alpha_{1}/2)\,\sigma(y,\alpha_{1};P_{0}),$
$m(y,\alpha_{1};\hat{P})+q\,\sigma(y,\alpha_{1};\hat{P})>m(y,\alpha_{1};P_{0})-\Phi^{-1}(\alpha_{1}/2)\,\sigma(y,\alpha_{1};P_{0})|Y_{T}=y],$
because $\Phi^{-1}(\alpha_{1}/2)<0$. For $\alpha_{2}\in(0,1)$, possibly different from $\alpha_{1}$, there exists a value $q_{0}(y,\alpha_{1};\alpha_{2})$ such that:
$\Pi_{0}[y,\alpha_{1};q_{0}(y,\alpha_{1};\alpha_{2})]=1-\alpha_{2}.$ (5.10)
Asymptotically, we get the $1-\alpha_{2}$ coverage probability of the true conditional prediction interval, although the true predictive density $P_{0}$ and the true distribution of $\hat{P}$ remain unknown.
Therefore, equations (5.9) and (5.10) can be replaced by their bootstrapped counterparts obtained from $S$ replicated paths of the series. More precisely, the bootstrapped conditional coverage probability is defined as:
@@ 5.3 Confidence Interval of the Prediction Interval
$\delta^{s}=\left\{\begin{array}[]{ll}1,&\mbox{if}\;\hat{m}^{s}(y,\alpha_{1})-q\,\hat{\sigma}^{s}(y,\alpha_{1})<\hat{m}(y,\alpha_{1})+\Phi^{-1}(\alpha_{1}/2)\,\hat{\sigma}(y,\alpha_{1}),\\ &\mbox{and}\;\hat{m}^{s}(y,\alpha_{1})+q\,\hat{\sigma}^{s}(y,\alpha_{1})>\hat{m}(y,\alpha_{1})-\Phi^{-1}(\alpha_{1}/2)\,\hat{\sigma}(y,\alpha_{1}),\\ 0,&\mbox{otherwise}.\end{array}\right.$
Then, we consider a solution $\hat{q}(y,\alpha_{1},\alpha_{2})$ of:
$\hat{\Pi}^{s}[y,\alpha_{1},\hat{q}(y,\alpha_{1},\alpha_{2})]=1-\alpha_{2}.$ (5.11)
@@
ensuring the $1-\alpha_{2}$ coverage probability. The bootstrap confidence set of the prediction interval is:
$\widehat{CSPI}(y,\alpha_{1},\alpha_{2})=\left\{m(y,\alpha_{1},\hat{P})\pm\hat{q}(y,\alpha_{1},\alpha_{2})\sigma(y,\alpha_{1},\hat{P})\right\}.$ (5.12)
This bootstrap confidence set is such that:
$\lim_{T\rightarrow\infty}\lim_{S\rightarrow\infty}P_{0}[\widehat{CSPI}(y,\alpha_{1},\alpha_{2})\supset PI(y,\alpha_{1})|Y_{T}=y]=1-\alpha_{2},\;\forall P_{0}.$ (5.13)
@@
Hence, the length of the estimated $\widehat{PI}(y,\alpha_{1})$ is enlarged by a factor $\hat{q}(y,\alpha_{1},\alpha_{2})/|\Phi^{-1}(\alpha_{1}/2)|$, that depends on the observed value $y$, in general. In practice, we can choose $\alpha_{1}=\alpha_{2}=0.05$ corresponding to the standard levels for prediction and confidence intervals, respectively. One could also choose $\alpha_{1}$ different from $\alpha_{2}$, including $\alpha_{1}=1.0$, which would correspond to the confidence set of point prediction equal to the median of the predictive density.

### v4

@@ 4.2.1 True and Estimated Prediction Intervals
where $P_{0}$ is the true density function of process ($Y_{t}$) and $Q(y,\alpha,P_{0})$ denotes the $\alpha$-quantile of $Y_{1,T+1}$ conditional on $Y_{T}=y$, derived from the joint multivariate predictive density $l(y|\underline{Y}_{T})$ (see Proposition 1 for the closed-form expression of the predictive density). Let $Q_{l}$ and $Q_{u}$ denote the true $\alpha_{1}/2$ and $21-\alpha_{1}/$ 2 conditional quantiles, respectively. Recall that under the semi-parametric approach, the true density $P_{0}$ is characterized by the parameter matrix $\Phi_{1,0}$ and functional parameter $g_{0}$. Then, conditional on the information at time $T$, the prediction interval is $PI(Y_{T},\alpha_{1})$. This is a random interval (set), a function of $Y_{T}$, and its conditional coverage level is $P_{0}(Y_{1,T+1}\in PI(Y_{T},\alpha_{1})|Y_{T})=1-\alpha_{1},\forall Y_{T}.$
Example 1: Linear AR(1) model, cont.: The theoretical prediction interval is: $PI(y,\alpha_{1})=(\phi y-Q_{0}(\alpha_{1}/2),\phi y+Q_{0}(1-\alpha_{1}/2)),$ where $Q_{0}$ is the quantile function corresponding to the true distribution function of $\epsilon_{t}$.
In the mixed VAR model $\epsilon_{t}$ is no longer a causal innovation and the marginal quantiles $Q_{0}$ have to be replaced by the conditional quantiles $Q(y,\alpha_{1},P_{0})$ depending on the current value $Y_{T}=y$. Moreover, this conditional quantile is not an affine function of $y$.
By using the standard expression of a Gaussian prediction interval with $\Phi$ denoting the Gaussian cumulative distribution function (c.d.f), the asymptotically valid prediction interval (4.1) for $Y_{1,T+1}$ can be equivalently written as:
@@
with [^5] Note that $\Phi^{-1}(\alpha_{1}/2)$ is negative. $m(y,\alpha_{1};P_{0})=0.5[Q_{l}(y,\alpha_{1}P_{0})+Q_{u}(y,\alpha_{1};P_{0})]$ and $\sigma(y,\alpha_{1};P_{0})=-[1/(2\Phi^{-1}(\alpha_{1}/2)]$
@@
$[Q_{u}(y,\alpha_{1};P_{0})-Q_{l}(y,\alpha_{1};P_{0})]$. This normalized representation of prediction interval resembling the traditional Gaussian approach can be used even if the conditional density function of $(Y_{t})$ is not Gaussian. In particular, the functions $m(y,\alpha_{1};P_{0})$ and $\sigma(y,\alpha_{1};P_{0})$ are nonlinear in $y$, in general.
@@
ii) Estimated Prediction Interval: The unknown marginal predictive density function $P_{0}$ can be consistently estimated from eq. (3.1) and denoted by $\hat{P}$, given the estimated matrix of autoregressive parameters $\hat{\Phi}_{1}$ and the residuals $\hat{\varepsilon}_{t}$ obtained from the semi-parametric GCov estimator, along with the nonparametric estimator $\hat{g}$. Then, the estimated prediction interval for $Y_{1,T+1}$ is:
$\widehat{PI}(y,\alpha_{1})=[m(y,\alpha_{1};\hat{P})\pm\Phi^{-1}(\alpha_{1}/2)\,\sigma(y,\alpha_{1};\hat{P})]=[Q_{l}(y,\alpha_{1};\hat{P}),Q_{u}(y,\alpha_{1};\hat{P})],$ (4.3)
where $Q_{l}(y,\alpha_{1};\hat{P})$ and $Q_{u}(y,\alpha_{1};\hat{P})$ are the $\alpha_{1}/2$ and $21-\alpha_{1}/$ 2 conditional quantiles of the estimated predictive density. This estimated prediction interval (4.3) is consistent of the true prediction interval (4.2) when the number of observations tends to infinity. We know that the estimated prediction interval $\widehat{PI}(Y_{T},\alpha_{1})$ does not always satisfy the conditional coverage condition in finite sample: $P_{0}[Y_{1,T+1}\in\widehat{PI}(Y_{T},\alpha_{1})|Y_{T}]=1-\alpha_{1},\;\;\forall Y_{T}$. In general, we have: $P_{0}(Y_{1,T+1}\in\widehat{PI}(Y_{T},\alpha_{1})|Y_{T})=1-\alpha_{1}(Y_{T}),$ i.e. the conditional coverage depends on $Y_{T}$, and the unconditional coverage: $P_{0}(Y_{1,T+1}\in\widehat{PI}(Y_{T},\alpha_{1}))=E_{0}[1-\alpha_{1}(Y_{T})],$ differs from $1-\alpha_{1}$ [^6] Because the process is Markov of order 1, we condition on $Y_{T}$ only instead of $\underline{Y}_{T}=(Y_{T},Y_{t-1},...)$, even though the estimators depend on the entire past..
Example 1: Linear AR(1) model, cont.: The estimated prediction interval becomes: $\widehat{PI}(Y_{T},\alpha_{1})=[\hat{\phi}Y_{T}-\hat{Q}(\alpha_{1}/2),\hat{\phi}Y_{T}+\hat{Q}(1-\alpha_{1}/2)].$ In the linear dynamic model, the estimated quantiles are easily derived from the residuals $\hat{\epsilon}_{t}=y_{t}-\hat{\phi}y_{t-1}$ ranked in increasing order.
Such a simple non-parametric estimation method of the quantile function is not available in the nonlinear dynamic framework, including the mixed VAR(1) model, where $\epsilon_{t}$ is no longer a causal innovation. Instead, the conditional quantiles are estimated from the kernel functional estimator of the predictive density. Note also that the estimates of $Q$ are expected to be less accurate than the estimate of the autoregressive coefficient $\phi$, because they approximate non-parametrically the tails of a multivariate distribution.
iii) 4.2.2 Backward Bootstrap Adjusted Prediction Interval
@@
Since the estimated prediction interval is random, its finite sample distribution can be approximated by a ”backward” bootstrap, i.e. by replicating the trajectory of the process by backcasting, conditional on $Y_{T}$ to adjust for the bias in coverage conditional on $Y_{T}$. More precisely, given $\hat{\Phi}_{1}$, $\hat{g}$ considered fixed and the residuals, we can generate by backcasting the artificial paths $Y_{t}^{s},\;t=1,...,T$, with the same terminal condition $Y_{T}^{s}=Y_{T}=y$ for all the bootstrapped samples (see, Corollary 2 for the closed-form expression of the backward predictive distribution). We propose the following algorithm:
@@
step 2: By replicating the backcasted path $S$ times, we end up generating $S$ bootstrapped series $Y_{t}^{s},t=1,...,T$, of length $T$ equal to the length of the initial series and with the same terminal value $Y_{T}$.
step 3: From each replicated path $(Y_{t}^{s},t=1,...,T)$, we estimate the model parameters $\Phi^{s}$ and $\hat{g}^{s}$, $s=1,...,S$. This allows us for computing at $Y_{T}$ the predictive density estimator $\hat{P}^{s},s=1,...,S$ of $Y_{T+1}$ given $Y_{T}$ and $S$ new prediction intervals for $Y_{1,T+1}$ from each of the replicated paths.
@@
where $\hat{P}^{s}$ is the semi-parametric estimate of $P_{0}$ from the generated path $(Y_{t}^{s},t=1,...,T)$. This bootstrap PI of $Y_{1,T+1}$ given $Y_{T}$ can be replicated independently $S$ times. The components of the prediction interval (4.4) are denoted by:
@@
step 5: For large $S$, the joint sample distribution of $[\hat{m}^{s}(y,\alpha_{1}),\hat{\sigma}^{s}(y,\alpha_{1})]$ provides an approximation of the conditional distribution of $[m(y,\hat{P}),\sigma(y,\hat{P})]$ given $Y_{T}$, when $T$ is finite and sufficiently large.
Example 1: Linear AR(1) model, cont.: In the linear AR(1) model, the bootstrap is usually applied by drawing independently in the sample distribution of residuals. This is justified by the additive decomposition of the conditional quantile.
In a nonlinear dynamic model with nonlinear dependence between $y$ and $Q$, the bootstrap has to be performed conditional on $Y_{T}$ to adjust for the bias in coverage conditional on $Y_{T}$.
4.3 Confidence Set of the Prediction Interval
The estimated prediction interval $\widehat{PI}(y,\alpha_{1})$ is a pointwise estimator of interval $PI(y,\alpha_{1})$ and, as any estimator, is random itself. This randomness is difficult to assess since $\widehat{PI}(y,\alpha_{1})$ is a random interval [see Molchanov and Molinari (2018) for random sets in Econometrics]. Let us now extend the method of pointwise estimation to build confidence sets for $PI(y,\alpha_{1})$. Since there does not exist a total ordering on intervals, we constrain the confidence set to be also of the form:
$\widehat{PI}(y,\alpha_{1},q)=[m(y,\alpha_{1};\hat{P})\pm q\,\sigma(y,\alpha_{1};\hat{P})],$ (4.6)
and search for an estimator of $q$ providing the correct asymptotic coverage. This confidence set has the following conditional coverage probability of the true prediction interval:
$\begin{array}[]{lcl}\Pi_{0}(y,\alpha_{1};q)&=&P_{0}[\widehat{PI}(y,\alpha_{1},q)\supset PI(y,\alpha_{1})|Y_{T}=y]\\ &=&P_{0}[m(y,\alpha_{1};\hat{P})-q\,\sigma(y,\alpha_{1};\hat{P})<m(y,\alpha_{1};P_{0})+\Phi^{-1}(\alpha_{1}/2)\,\sigma(y,\alpha_{1};P_{0}),\\ m(y,\alpha_{1};\hat{P})&+&q\,\sigma(y,\alpha_{1};\hat{P})>m(y,\alpha_{1};P_{0})-\Phi^{-1}(\alpha_{1}/2)\,\sigma(y,\alpha_{1};P_{0})|Y_{T}=y]\hfill(4.7)\end{array}$
because $\Phi^{-1}(\alpha_{1}/2)<0$. For $\alpha_{2}\in(0,1)$, possibly different from and less than $\alpha_{1}$, there exists a value $q_{0}(y,\alpha_{1};\alpha_{2})$ such that:
$\Pi_{0}[y,\alpha_{1};q_{0}(y,\alpha_{1};\alpha_{2})]=1-\alpha_{2},\;\forall y.$ (4.8)
Asymptotically, we get the conditional $1-\alpha_{2}$ coverage probability of the true conditional prediction interval, although the true predictive density $P_{0}$ and the true distribution of $\hat{P}$ remain unknown. In practice, this procedure can provide us a prediction interval at level 95%, for example, when the available sample does not contain sufficiently many tail observations and a reliable prediction interval at level 90% or lower is only estimable.
Then, equations (4.7) and (4.8) can be replaced by their bootstrapped counterparts obtained from $S$ replicated paths of the series, which are backcast prior to T, given $Y_{T}=y$. Next, a forecast at T+1 is computed from each of the replicated paths. More precisely, the bootstrapped conditional coverage probability is defined as:
@@ 4.3 Confidence Set of the Prediction Interval
where $\delta^{s}=\left\{\begin{array}[]{ll}1,&\mbox{if}\;\hat{m}^{s}(y,\alpha_{1})-q\,\hat{\sigma}^{s}(y,\alpha_{1})<\hat{m}(y,\alpha_{1})+\Phi^{-1}(\alpha_{1}/2)\,\hat{\sigma}(y,\alpha_{1}),\\ &\mbox{and}\;\hat{m}^{s}(y,\alpha_{1})+q\,\hat{\sigma}^{s}(y,\alpha_{1})>\hat{m}(y,\alpha_{1})-\Phi^{-1}(\alpha_{1}/2)\,\hat{\sigma}(y,\alpha_{1}),\\ 0,&\mbox{otherwise}.\end{array}\right.$
Then, we consider a solution $\hat{q}^{S}(y,\alpha_{1},\alpha_{2})$ of:
$\hat{\Pi}^{s}[y,\alpha_{1},\hat{q}^{S}(y,\alpha_{1},\alpha_{2})]=1-\alpha_{2}.$ (4.9)
@@
ensuring a $1-\alpha_{2}$ conditional coverage probability. The bootstrap confidence set for the prediction interval is:
$\widehat{CSPI}(y,\alpha_{1},\alpha_{2})=\left\{m(y,\alpha_{1},\hat{P})\pm\hat{q}^{S}(y,\alpha_{1},\alpha_{2})\sigma(y,\alpha_{1},\hat{P})\right\},$ (4.10)
@@
and $\lim_{T\rightarrow\infty}\lim_{S\rightarrow\infty}P_{0}[\widehat{CSPI}(y,\alpha_{1},\alpha_{2})\supset PI(y,\alpha_{1})$ $|Y_{T}=y]=1-\alpha_{2},\;\forall y,\forall P_{0}.$ Hence, the length of the estimated $\widehat{PI}(y,\alpha_{1})$ is modified by a factor $\hat{q}^{S}(y,\alpha_{1},\alpha_{2})/|\Phi^{-1}(\alpha_{1}/2)|$, that depends on the observed value $Y_{T}=y$, in general. In practice, we can choose $\alpha_{1}=\alpha_{2}=0.05$ corresponding to the standard levels for prediction and confidence intervals, respectively. As mentioned earlier, we can choose $\alpha_{1}$ different from $\alpha_{2}$, including $\alpha_{1}=1.0$, which would correspond to the confidence set for a point prediction equal to the median of the predictive density.